# BESSTIE: A Benchmark for Sentiment and Sarcasm Classification for Varieties of English

### v1

@@ 4.3. Data Quality & Statistics
With respect to the true label (based on the location or topic), results show higher agreement with the en-IN annotator ($\kappa$ = 0.41) compared to the en-UK annotator ($\kappa$ = 0.34), indicating fair to moderate consistency. The inter-annotator agreement itself is $0.26$. Figure 2 shows that annotators find it difficult to agree on inner-circle varieties, and have the highest agreement when identifying en-IN variety, with 46 agreements. While agreement between annotators is limited for inner-circle varieties, the higher agreement for the en-IN indicates that the subset reliably represents this variety, thereby validating our collection methodology. As a result, the two subsets (reviews and comments) collectively form a good representative sample for language varieties.
Note: We also attempt to fine-tune an encoder model fine-tuned on an unlabeled dialectal corpus, to predict the language varieties. However, this was unsuccessful due to: (a) unavailability of the dataset for en-UK, (b)the dataset being dated and inaccurate in the domain of the collected data.
@@
Table 2. Reviews and comments from different varieties with their corresponding annotated sentiment and sarcasm labels.
Dataset Statistics
The benchmark dataset contains annotations for two tasks: sentiment and sarcasm classification, distributed across the three varieties. Table 3 provides a breakdown of the annotated dataset (consisting of google and reddit), including the number of samples in each split (Train, Validation, and Test), along with average word length and language probability. The % Positive sentiment and % Positive sarcasm indicate the proportion of samples with positive sentiment and true sarcasm respectively. These values are the same for train, validation and test set since we perform stratified sampling. We present a few example comments along with their corresponding sentiment and sarcasm labels in Table 2. The examples illustrate the diversity of language use and the challenges involved in detecting both sentiment and sarcasm across the varieties.
Variety Subset Train Valid Test % Positive Sentiment % Positive Sarcasm $P(\bar{\texttt{eng}})$ $\bar{w}$
en-AU google 946 130 270 73% 7% 0.99 63.97
reddit 1763 241 501 32% 42% 0.98 51.72
en-IN google 1648 225 469 75% 1% 0.99 44.34
reddit 1686 230 479 25% 13% 0.87 26.92
en-UK google 1817 248 517 75% 0% 0.99 72.21
reddit 1007 138 287 12% 22% 0.98 38.04
Table 3. Dataset statistics for each variety, subset, split, and label type. $\mu$ represents the average probability of English. $P(\bar{\texttt{eng}})$ is the average language probability (calculated using fastText (Grave et al., 2018) word vectors) of a review or comment being in English. $\bar{w}$ denotes average number of words in a review/comment.
5. Evaluation Methodology
With the annotated dataset in place, we now discuss the evaluation methodology used to evaluate the LLMs on the dataset.
The training strategies for encoder and decoder models are as follows. We train encoder models using cross-entropy loss, weighted on class distribution:
$\mathcal{L}_{\texttt{{Encoder}}}=-\frac{1}{N}\sum_{j=1}^{N}\sum_{k=1}^{C}w_{k}\cdot y_{j,k}\cdot\log p_{j,k}$
Here, $N$ denotes the total number of samples in the batch, $C$ represents the number of classes (2 since the tasks are binary classification), and $w_{k}$ is the weight assigned to class k, calculated based on the class distribution ($w_{\texttt{neg}}$ for the negative and $w_{\texttt{pos}}$ for the positive class). $y_{j}$ is the expected output label for the $j$th instance where $y_{j,1}=1$ if the label is 1 and $y_{j,0}=1$ if the label is 0. $p_{j,k}$ is the probability of the $j$th instance for $k$th class.
In contrast, for decoder models, we perform zero-shot instruction fine-tuning (Ouyang et al., 2022), using maximum likelihood estimation – a standard learning objective for causal language modeling (Jain et al., 2023):
$\mathcal{L}_{\texttt{{Decoder}}}=-\frac{1}{N}\sum_{j=1}^{N}\left\{\sum_{i=|\mathbf{x}^{j}|+1}^{|\mathbf{x}^{j}|+|\mathbf{t}^{j}|}\log p(\mathbf{x}_{i}^{j}|\mathbf{x}_{<i}^{j})\right\}$
Here, $\mathbf{x}_{<i}^{j}=[\mathbf{x}_{1}^{j},\ldots,\mathbf{x}_{i-1}^{j}]$ denotes the subsequence before $\mathbf{x}_{i}^{j}$ and ||$\cdot$|| ⋅ | is the number of tokens. Specifically, $|t_{j}|$ is the number of generated tokens and is set to $1$. In other words, given the text and a task-specific prompt (given in Table 4555The forms are experimentally determined using a few test examples.), the expected behaviour of the LLM is to generate either 1 (for positive) or 0 (for negative).
Task Prompt
Sentiment Classification “Generate the sentiment of the given text. Do not give an explanation.”
Sarcasm Classification “Predict if the given text is sarcastic. Do not give an explanation.”
Table 4. Task-specific prompts used to fine-tune decoder models. Both prompts restrict the model from generating an explanation.
6. Experiment Details
We conduct our experiments on a total of nine LLMs. These include six encoder models, three trained on English corpora: BERT-Large (bert(Devlin et al., 2019), RoBERTa-Large (roberta(Liu et al., 2019), ALBERT-XXL-v2 (albert(Lan et al., 2020); and three multilingual models: Multilingual-BERT-Base (mbert), Multilingual-DistilBERT-Base (mdistil(Sanh et al., 2020), XLM-RoBERTa-Large (xlm-r(Goyal et al., 2021). We also evaluate three decoder models, one trained on English corpora: Gemma2-27B-Instruct (gemma(Gemma Team et al., 2024) and two multilingual models: Mistral-Small-Instruct-2409 (mistral(Jiang et al., 2023), Qwen2.5-72B-Instruct (qwen(Yang et al., 2024). The model-specific details including architecture, language coverage of pre-training corpus, and number of parameters are described in Table 5.
While encoder models are fine-tuned with full precision, we fine-tune quantized decoder models using QLoRA (Dettmers et al., 2023) adapters, targeting all linear layers. All models are fined-tuned for 30 epochs, with a batch size of 8 and Adam optimiser. We choose an optimal learning rate by performing a grid search over the following values: 1e-5, 2e-5, and 3e-5. All experiments are performed using two NVIDIA A100 80GB GPUs.
We report our task performances on three macro-averaged F-Score, i.e., unweighted average, disregarding the imbalance in label distribution.
@@ 6. Experiment Details
Table 5. Details of models including architecture (Arch.), language coverage in pretraining data (Lang.), and number of trainable parameters (# Params). Here, M is millions and B is billion.
7. Results
Figure 3. Performance comparison of various models on the sentiment classification task across different English varieties (en-AU, en-IN, and en-UK).
7.1. Overall results
We first present our results on the two tasks where models are trained and evaluated on the same variety. Figure 3 and 4 describe the model performances, reported using F-Score, on the sentiment and sarcasm classification tasks respectively. While sentiment classification is performed using both google and reddit subsets, due to the absence of sarcasm labels in google subset (as shown in Table 3), sarcasm classification is conducted only on reddit subset.
Sentiment Classification
Figure 4. Performance comparison of various models on the sarcasm classification task across different English varieties (en-AU, en-IN, and en-UK).
Sarcasm Classification
Similarly, Table 7 shows that, for sarcasm classification, both mistral and mbert achieve the highest average performance across all varieties, with an F-Score of 0.68. In contrast, gemma again reports the lowest average performance across all varieties, with an average F-Score of 0.45. The lower performance on sarcasm classification as compared to sentiment classification is expected because sarcasm often relies on not just subtle linguistic cues but also localised contemporary contextual information (Abercrombie and Hovy, 2016).
Model Enco. Mono. google-Sentiment reddit-Sentiment reddit-Sarcasm
albert 0.88 0.75 0.65
bert 0.88 0.80 0.60
roberta 0.89 0.79 0.55
mbert 0.88 0.65 0.68
mdistil 0.86 0.78 0.51
xlm-r 0.77 0.73 0.62
gemma 0.79 0.60 0.45
mistral 0.91 0.84 0.68
qwen 0.45 0.82 0.52
[…]
Table 6. Model performances, averaged across all varieties, on sentiment and sarcasm classification tasks across two subsets (google and reddit). $\mu$ represents the average performance across all models. ‘Enco.’ represents encoder models. ‘Mono.’ represents monolingual models.
7.2. Impact of Domain and Models
The average results for the models highlight the need to further probe into factors influencing model performances, namely, domain, model properties and language varieties. The results in Tables 7 and 8 describe the model performances (reporting F-Score), grouped based on model properties. Finally, task performances, averaged across all models, on varieties are presented in Table 9.
google vs. reddit
Our findings (the last row $\mu$ in Table 6), indicate that models perform better on google subset, achieving an F-Score of 0.81, compared to an F-Score of 0.75 on reddit subset. This difference can be explained by the distinct writing styles of the two sources. Reviews in google are generally more formal and informative, while posts and comments from reddit often exhibit a short (as shown in Table 3), conversational tone typical of social media. Additionally, posts and comments in reddit come from forums frequented by local speakers, which means the language and expressions used are more reflective of the variety and cultural nuances.
Domain-Task Enco. Deco.
google-Sentiment 0.86 0.72
reddit-Sentiment 0.75 0.75
reddit-Sarcasm 0.60 0.55
[…]
Table 7. Task performances, averaged across all varieties, grouped based on model architectures. $\mu$ represents the average performance across all domain-task pairs.‘Enco.’ represents encoder models while ‘Deco.’ represents decoder models.
Domain-Task Mono. Mult.
google-Sentiment 0.86 0.78
reddit-Sentiment 0.73 0.76
reddit-Sarcasm 0.56 0.60
[…]
Table 8. Task performances, averaged across all varieties, grouped based on language diversity in pre-training data. ‘Mono.’ represents monolingual models, and ‘Mult.’ represents multilingual models. $\mu$ represents the average performance across all domain-task pairs.
Model properties
Table 7 shows a consistent trend where encoder models report higher performance than decoder models across different tasks. This performance gap is expected as encoder architecture is inherently better suited for sequence classification tasks, while decoder architecture is optimised for text generation tasks. Table 8 shows that, although marginal, monolingual models report a higher average performance compared to multilingual models. This performance gap is prominent on the task involving google subset, while multilingual models perform better on tasks involving reddit subset.
Domain-Task en-AU en-IN en-UK
google-Sentiment 0.94 0.64 0.86
reddit-Sentiment 0.78 0.69 0.78
reddit-Sarcasm 0.62 0.56 0.58
[…]
Table 9. Task performances, averaged across all models, on all varieties. $\mu$ represents the average performance across all domain-task pairs.
Comparison of language varieties
Finally, Table 9 shows that models, on average, perform better on both tasks when trained on en-AU variety, closely followed by performances on en-UK variety. Both en-AU and en-UK are inner-circle varieties. Conversely, models consistently perform the worst on en-IN, an outer-circle variety.
Figure 5. Cross-variety performance analysis of mistral. The figure compares three different scenarios: pre-trained (PT), in-variety fine-tuning, and cross-variety fine-tuning for sentiment and sarcasm classification across all varieties.
Figure 6. Cross-variety performance analysis of bert. The figure compares two different scenarios: in-variety fine-tuning, and cross-variety fine-tuning for sentiment and sarcasm classification across all varieties.
Figure 7. Cross-domain performance analysis of mistral and bert comparing a decoder and an encoder model.
7.3. Cross-variety/domain evaluation
Finally, we perform cross-variety/cross-domain evaluation. This refers to the scenario where the datasets used to train and test a model are from different groups (i.e., trained on en-AU, tested on en-IN, for example). The following discussion is focused on generalisation: “can models trained on one language variety generalise well to others”.
Cross-variety evaluation
As seen earlier, the model performs significantly worse on sarcasm classification compared to sentiment classification. While in-variety evaluation shows that fine-tuning improves over the pre-trained model with substantial increases in F-Score, it negatively impacts cross-variety generalization. This means there is a notable decrease in F-Score across other varieties when the model is fine-tuned on a specific variety. This supports the idea that sarcasm classification requires the model to grasp linguistic and cultural nuances specific to a variety, making generalisation across varieties challenging. The corresponding results for bert are in Figure 6.
Cross-domain evaluation

### v3

@@ 2.2 Quality Assessment
With respect to the true label (based on the location or topic), results show higher agreement with the en-IN annotator ($\kappa$ = 0.41) compared to the en-UK annotator ($\kappa$ = 0.34), indicating fair to moderate consistency. The inter-annotator agreement itself is $0.26$. Figure 2 shows that annotators find it difficult to agree on an inner-circle variety, and have the highest agreement when identifying the en-IN variety, with 46 agreements, demonstrating that the subset reliably represents this variety. As a result, the two subsets (reviews and comments) collectively form a good representative sample for language varieties.
Variety Subset $P(\bar{\texttt{eng}})$ $P(\bar{\texttt{v}})$ F-Score en-AU google 0.99 0.99 0.99 reddit 0.98 0.95 0.93 en-IN google 0.99 0.94 0.91 reddit 0.87 0.78 0.69 en-UK google 0.99 0.99 0.99 reddit 0.98 0.93 0.90
Table 3: Data quality for each variety and subset. $P(\bar{\texttt{eng}})$ is the average language probability (calculated using fastText Grave et al. (2018) word vectors) of a review or comment being in English. $P(\bar{\texttt{v}})$ represents the average variety probability of a review or comment being in the corresponding variety. F-Score measures the performance of the variety predictor.
Automated validation
We use two predictors, a language predictor, and a variety predictor, to perform an automatic evaluation of our dataset. For language predictor, we use fastText word vectors and extract the probability of the text being in English. We then fine-tune distil on ICE-Corpora, to use as a variety predictor. We model the task as binary classification (i.e., inner-circle vs. outer-circle variety classification). We use the ICE-Australia Smith and Peters (2023) and ICE-India subsets of the corpus [^3] Due to lack of access to ICE-GB, we do not include en-UK variety in training the predictor., focusing on the transcriptions from unprompted monologues and private dialogues (S1A and S2A headers). Similar to language predictor, we extract the probability of text being in the corresponding variety.
$P(\bar{\texttt{eng}})$ in Table 3 is the probability that the text is written in English. $P(\bar{\texttt{v}})$ is the average variety probability. High values indicate that the variety predictor is capable of discerning text written in inner-circle varieties from those written in the outer-circle variety (i.e., en-IN). The lower language and variety probabilities on the reddit subset from en-IN, with 0.87 and 0.78 respectively, can be attributed to the code-mixed text in the subset. These results along with our manual annotation exercise show that the collected data represents the three language varieties. The probabilities from the model are in line with the ground truth, as measured using the F-Score.
Variety Sent. Sarc. en-AU 0.61 0.47 en-IN 0.65 0.51 en-UK 0.79 0.63
Table 4: Inter-annotator agreement for the validation annotation exercise. Agreement is measured between the original annotator and an independent annotator for each language variety. Here Sent. and Sarc. denote sentiment and sarcasm labels respectively.
Variety Subset Train Valid Test % Pos. Sent. % Pos. Sarc. Avg. no. of words en-AU google 946 130 270 73% 7% 63.97 reddit 1763 241 501 32% 42% 51.72 en-IN google 1648 225 469 75% 1% 44.34 reddit 1686 230 479 25% 13% 26.92 en-UK google 1817 248 517 75% 0% 72.21 reddit 1007 138 287 12% 22% 38.04
Table 5: Dataset statistics for each variety, subset, split, and label type. % Pos. Sent. and % Pos. Sarc. indicates the proportion of samples with positive sentiment and true sarcasm respectively.
2.3 Annotation
We hire one annotator [^4] Two of these are the authors of the paper who were also involved in the evaluation. each for the three language varieties. The annotators assign the processed reviews and posts two labels: sentiment and sarcasm. The choice of labels given are negative, positive, and discard. We instruct the annotators to use the discard label for uninformative examples with no apparent polarity and in a few cases, for computer-generated messages. The texts with discard label are discarded.
We conduct an additional annotation exercise to evaluate the reliability of the sentiment and sarcasm annotations. For this, we randomly selected 50 instances from each variety, annotated by the original annotators, and hire a new independent annotator for each variety. The inter-annotator agreement between in the original and the independent annotators, measured by $\kappa$, are reported in the Table 4. The absolute agreement values for sarcasm, though lower than those for sentiment, are comparable to prior work. In particular, Joshi et al. (2016) report a $\kappa$ of 0.44 on their sarcasm labels. These scores suggest that both sentiment and sarcasm labels in our dataset are annotated with high degree of reliability. We compensate the annotators at the casual employment rate, equivalent to 22 USD per hour, as prescribed by the host institution. We provide the annotation guideline in Appendix A.
2.4 Dataset Statistics
The resultant besstie dataset contains annotations for two tasks: sentiment and sarcasm classification, for the two domains and three varieties. Table 5 provides a breakdown of the annotated dataset (consisting of google and reddit), including the number of samples in each split (Train, Validation, and Test), along with average word length. The % Pos. Sent. and % Pos. Sarc. indicates the proportion of samples with positive sentiment and true sarcasm respectively. These values are the same for the train, validation, and test set since we perform stratified sampling. We present example comments along with their corresponding sentiment and sarcasm labels in Appendix B. We also confirm that the text collected for Reddit subset of the besstie is generated in 2024. While this subset represents a static snapshot, its recency allows it to reflect contemporary linguistic usage effectively.
Figure 3: Performance comparison of various models on the sentiment classification task across different English varieties (en-AU, en-IN, and en-UK).
Figure 4: Performance comparison of various models on the sarcasm classification task across different English varieties (en-AU, en-IN, and en-UK).
3 Experiment Details
We conduct our experiments on a total of nine LLMs. These include six encoder models, three pre-trained on English corpora: BERT-Large (bertDevlin et al. (2019), RoBERTa-Large (robertaLiu et al. (2019), ALBERT-XXL-v2 (albertLan et al. (2020); and three multilingual models: Multilingual-BERT-Base (mbert), Multilingual-DistilBERT-Base (mdistilSanh et al. (2020), XLM-RoBERTa-Large (xlm-rGoyal et al. (2021). We also evaluate three decoder models, one pre-trained on English corpora: Gemma2-27B-Instruct (gemmaGemma Team et al. (2024) and two multilingual models: Mistral-Small-Instruct-2409 (mistralJiang et al. (2023), Qwen2.5-72B-Instruct (qwenYang et al. (2024). The model-specific details including architecture, language coverage of pre-training corpus, and number of parameters are described in Appendix C.
While encoder models are fine-tuned with full precision, we fine-tune quantised decoder models using QLoRA Dettmers et al. (2023) adapters, targeting all linear layers. All models (including the variety predictor) are fined-tuned for 30 epochs, with a batch size of 8 and Adam optimiser. We choose an optimal learning rate by performing a grid search over the following values: 1e-5, 2e-5, and 3e-5. All experiments are performed using two NVIDIA A100 80GB GPUs.
We fine-tune encoder models using cross-entropy loss, weighted on class distribution (positive: ‘1’ and negative: ‘0’ labels). In contrast, for decoder models, we perform zero-shot instruction fine-tuning Ouyang et al. (2022), using maximum likelihood estimation – a standard learning objective for causal language modeling Jain et al. (2023).
For decoder models, we use two task-specific prompts, one for each task. For sentiment classification, we the prompt the model with:
“Generate the sentiment of the given text. 1 for positive sentiment, and 0 for negative sentiment. Do not give an explanation.”
Similarly, for sarcasm classification, we use:
“Predict if the given text is sarcastic. 1 if the text is sarcastic, and 0 if the text is not sarcastic. Do not give an explanation.”
Given the text and a task-specific prompt, the expected behaviour of the LLM is to generate either 1 (for positive) or 0 (for negative). The forms of these prompts are experimentally determined using a few test examples. Finally, we report our task performances on three macro-averaged F-Score, i.e., unweighted average, disregarding the imbalance in label distribution.
4 Results
We present our results and subsequent discussions to address the following questions: (a) How well do current LLMs perform on the benchmark tasks? (Section 4.1); (b) How do factors such as model properties and domain affect the task performance? (Section 4.2); (c) Can models trained on one language variety generalise well to others? (Section 4.3).
4.1 Task Performance
We first present our results on the two tasks where models are trained and evaluated on the same variety. Figures 3 and 4 describe the model performances, reported using F-Score, on the sentiment and sarcasm classification tasks respectively. While sentiment classification is performed using both google and reddit subsets, due to the absence of sarcasm labels in google subset (as shown in Table 5), sarcasm classification is conducted only on reddit subset. Table 6 shows that models, on average, perform better on both tasks when trained on en-AU variety, closely followed by performances on en-UK variety. Conversely, models consistently perform the worst on en-IN.
Domain-Task en-AU en-IN en-UK
google-Sentiment 0.94 0.64 0.86
reddit-Sentiment 0.78 0.69 0.78
reddit-Sarcasm 0.62 0.56 0.58
[…]
Table 6: Task performances, averaged across all models, on all varieties. $\mu$ represents the average performance across all domain-task pairs.
Sentiment Classification
Sarcasm Classification
Similarly, for sarcasm classification, both mistral and mbert achieve the highest average performance across all varieties, with an F-Score of 0.68, as shown in Table 7. In contrast, gemma again reports the lowest average performance across all varieties, with an average F-Score of 0.45. The lower performance on sarcasm classification as compared to sentiment classification is expected because sarcasm often relies on not just subtle linguistic cues but also localised contemporary contextual information Abercrombie and Hovy (2016).
4.2 Impact of Models and Domain
The average results for the models highlight the need to further probe into factors influencing model performances, namely, model properties and domain.
Model properties
The results ($\mu$ values) in the corresponding columns of Table 7 describe the model performances (reporting F-Score), grouped based on model properties (encoder/decoder, mono/multilingual). The values show a consistent trend where encoder models report higher performance than decoder models across different tasks. This performance gap is expected as encoder architecture is inherently better suited for sequence classification tasks, while decoder architecture is optimised for text generation tasks. We also find that, although marginal, monolingual models report a higher average performance compared to multilingual models. This performance gap is prominent on the task involving google subset, while multilingual models perform better on tasks involving reddit subset.
google vs. reddit
Our findings (the last row $\mu$ for the task columns in Table 7), indicate that models perform better on google subset, achieving an F-Score of 0.81, compared to an F-Score of 0.75 on reddit subset. This difference can be explained by the distinct writing styles of the two sources. Reviews in google are generally more formal and informative, while posts and comments from reddit often exhibit a short (as shown in Table 5), conversational tone typical of social media. Additionally, posts and comments in the reddit subset come from forums frequented by local speakers, which means the language and expressions used are more reflective of the variety and cultural nuances. We present results from cross-domain evaluation in the Appendix D.
[…]
Table 7: Model performances, averaged across all varieties, on sentiment and sarcasm classification tasks across two subsets (google and reddit). $\mu$ represents the average performance across all models. ‘Enco.’ represents encoder models. ‘Mono.’ represents monolingual models.
Figure 5: Cross-variety performance analysis of mistral. The figure compares three different scenarios: pre-trained (PT), in-variety fine-tuning, and cross-variety fine-tuning for sentiment and sarcasm classification across all varieties.
4.3 Cross-variety Evaluation
Finally, we perform cross-variety evaluation. This refers to the scenario where the datasets used to train and test a model are from different groups (i.e., trained on en-AU, tested on en-IN, for example).
The model performs significantly worse on sarcasm classification compared to sentiment classification. While in-variety evaluation shows that fine-tuning improves over the pre-trained model with substantial increases in F-Score, it negatively impacts cross-variety generalisation. This means there is a notable decrease in F-Score across other varieties when the model is fine-tuned on a specific variety. This supports the idea that sarcasm classification requires the model to grasp linguistic and cultural nuances specific to a variety, making generalisation across varieties challenging.
Variety Sample dial coll cont code en-AU 70 9 28 6 - en-IN 90 97 33 3 8 en-UK 53 7 15 5 -
Table 8: Counts of identified features in misclassified examples from mistral for each variety. dial denotes dialect features, coll denotes locale-specific colloquial expressions, cont denotes instances requiring additional context, and code denotes occurrences of code-mixed text.
5 Error Analysis
We randomly sample up to 30 examples misclassified by mistral, from each variety and domain-task configuration. We then manually analyse these misclassified examples to identify pervasive or obligatory dialect features, as defined by eWAVE Kortmann et al. (2020). A dialect feature is defined as any lexical or syntactic variation from standard English. In addition, we identify other features, including locale-specific colloquial expressions, instances that require additional context, and occurrences of code-mixed text. Table 8 summarises the sample sizes and the counts of these identified features for each variety. Detailed examples for each feature type are provided in Appendix E.
The en-IN variety exhibits a high count of dialect features (97) relative to its sample size (90), suggesting that regional variations significantly challenge model performance. Locale-specific colloquial expressions are prevalent in all varieties (en-AU: 28; en-IN: 33; en-UK: 15), highlighting the impact of regional lexicons on model performance. These findings motivate future research in sentiment and sarcasm classification to better accommodate such colloquial expressions for inner-circle varieties, and to additionally adapt to dialect features for outer-circle varieties of English.
6 Related work
NLP Benchmarks such as GLUE Wang et al. (2018), SuperGLUE Wang et al. (2019), and DynaBench Kiela et al. (2021) have played a pivotal role in evaluating LLMs on multiple NLP tasks, including sentiment classification. However, these benchmarks are limited in their ability to capture the linguistic nuances of non-standard language varieties. Efforts to address this include datasets for African languages and dialects Muhammad et al. (2023), Arabic dialects Elmadany et al. (2023), and CreoleVal Lent et al. (2024) which provides datasets for 28 creole languages. More recent progress includes DialectBench Faisal et al. (2024), which presents a collection of dialectal datasets and associated NLP models across 281 dialectal varieties for several tasks. However, the benchmark does not contain English dialectal datasets for sentiment and sarcasm classification.
Similarly, several sarcasm classification datasets for standard variety of English exist, consisting of reddit comments Khodak et al. (2018), amazon reviews Filatova (2012), and tweets Ptáček et al. (2014); Painter et al. (2022); Abercrombie and Hovy (2016); Oprea and Magdy (2020). While datasets like ArSarcasm-v2 Abu Farha et al. (2021) present annotations for sarcasm, sentiment, and the dialect of each tweet for Arabic, there is a notable absence of sarcasm classification datasets that account for non-standard varieties of English. Similar to their work, we also create a dataset that contains both sentiment and sarcasm labels. To the best of our knowledge, besstie is the first benchmark for sentiment and sarcasm classification specifically for varieties of English.
7 Conclusion
This paper presents besstie, a benchmark for sentiment and sarcasm classification across three varieties of English: Australian (en-AU), Indian (en-IN), and British (en-UK). Our evaluation spans nine LLMs, six encoders and three decoders, assessed on datasets collected from Google Places reviews (google) and Reddit comments (reddit). The models perform consistently better on en-AU and en-UK (i.e., inner-circle varieties) than on en-IN (i.e., the outer-circle variety) for both tasks. We verify the variety representation of collected data using: (a) manual annotation and (b) automated validation. We also perform an additional annotation exercise to verify the reliability of the annotations. Although models report high performance on sentiment classification (F-Score of 0.81 and 0.75, for google and reddit respectively), they struggle with detecting sarcasm (F-Score of 0.59 on reddit), indicating that sarcasm classification is still largely unsolved. This is supported by our cross-variety evaluation which reveals limited generalisation capability for sarcasm classification, where cultural and contextual understanding is crucial. Notably, monolingual models marginally outperform multilingual models, suggesting that language diversity in pre-training does not extend to varieties of a language. Our error analysis provides potential considerations to drive future research in sentiment and sarcasm classification to better accommodate colloquial expressions for inner-circle varieties, and dialect features for outer-circle varieties of English. In conclusion, besstie provides a resource to measure bias in LLMs towards non-standard English varieties for the two tasks.
Limitations
Our assumption of national varieties as representative forms of dialect is a simplification. Within each locale, there are significant regional, sociolectal, and generational differences. For example, Australian English spoken in Sydney may differ significantly from that in Perth. The language used in online communications evolves rapidly, driven by cultural trends, viral content, and changing societal norms. Phrases, slang, and expressions can gain popularity and fade quickly. A static dataset, such as ours, is hence limited in reflecting these continuous and dynamic changes in language use. While our efforts to remove non-English text were thorough, there may still be instances where language mixing or code-switching occurs, especially in the Indian English subset, affecting model performance. The annotation process was limited by the number of human annotators and their subjective interpretations. While we perform an additional annotation exercise to validate the annotation process, use of a single annotator can introduce individual biases. We also note that comments from Reddit may differ from other social media sites, such as Twitter/X, in terms of user demographics and discourse style. This potentially influences the nature of topic-based discussions captured in our dataset. Our reliance on Reddit data is due to cost-related constraints associated with accessing Twitter/X API.
Ethical Considerations
The project received ethics approval from the Human Research Ethics Committee at UNSW, Sydney (reference number: iRECS6514), the host organisation for this research.We treat the reviews as separate texts and do not aggregate them by user in any way. The Reddit comments and Google reviews used in this paper are from publicly available posts, accessible through official APIs. We adhere to the terms of service and policies outlined by these platforms while gathering data. We do not attempt to de-identify any users or collect aggregate information about individual users. The dataset, including sentiment and sarcasm annotations, is shared in a manner that preserves user privacy and abides by the rules and guidelines of the source platforms.
Acknowledgements
This research was funded by Google’s Research Scholar grant, awarded to Aditya Joshi and Diptesh Kanojia in 2024.
Appendix A Annotation Guidelines
Task Overview
Your task involves determining whether each Google review or Reddit comment expresses positive or negative sentiment and whether it uses sarcasm.
Annotation Instructions
For Sentiment Classification:
  • Positive Sentiment (Label 1): Annotate a comment as positive if it reflects favorable emotions such as happiness, satisfaction, agreement, or excitement.
  • Negative Sentiment (Label 0): Annotate a comment as negative if it conveys unfavorable emotions such as sadness, disappointment, frustration, or criticism.
  • Neither Positive nor Negative (Discarded, Label 2): discard the comment if it does not convey clear positive or negative sentiment (mark the same sarcasm label as 2).
For Sarcasm Classification:
  • Sarcastic Comment (Label 1): Annotate if the comment uses irony or mockery to express contempt or ridicule.
  • Non-Sarcastic Comment (Label 0): Annotate if the comment is straightforward and does not employ sarcasm or irony.
Additional Guidelines
Consider the tone and context of the comment when assigning sentiment or sarcasm labels. For sarcasm, look for indicators such as exaggerated language, contradictions, or unexpected statements. Base your decision on the sentiment explicitly expressed in the text, avoiding personal biases. It is acceptable to mark a comment as undecided or seek clarification if you cannot confidently assign a sentiment label.
Appendix B Dataset Examples
@@
Table 9: Reviews and comments from different varieties with their corresponding annotated sentiment and sarcasm labels.
The examples presented in Table 9 illustrate the diversity of language use and the challenges involved in detecting both sentiment and sarcasm across the varieties.
Appendix C Additional Model Details
We describe additional model details, including architecture, the language coverage in their pre-training data, and the number of trainable parameters in Table 10.
@@ Appendix C Additional Model Details
Table 10: Details of models including architecture (Arch.), language coverage in pre-training data (Lang.), and number of trainable parameters (# Params). Here, M is millions and B is billion.
Appendix D Cross-domain Evaluation